OpenStudy (anonymous):

in a golf tournament players compeated in pairs joe and mellon were partners joe got a score of +13 and mellon got a score of -14 what was their scores combined

jimthompson5910 (jim_thompson5910):

13+(-14) = 13 - 14 = -1

OpenStudy (saifoo.khan):

negative one.